OpenStudy (anonymous):

Can someone help me with this set of equations y=x^2+5x-3 y-x=2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

second equation can be written as \(y=x+2\) right?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

yeah thats right @misty1212

OpenStudy (misty1212):

ok then the top y must be \(x+2\) also

OpenStudy (misty1212):

that means \[x+2=x^2+5x-3\]

OpenStudy (anonymous):

So substitute it... right

OpenStudy (misty1212):

yeah exactly now you have a quadratic equation to solve

OpenStudy (misty1212):

set it equal to zero by subtracting x and 2 from both sides then if you are lucky it will factor you good from there?

OpenStudy (anonymous):

I think so yes @misty1212 thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]